### 11 Mar 2026 v2

# On automatic boundedness of operators between ordered and topological vector spaces

Abstract
We study topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces. Levi and Lebesgue operators are investigated.
E. Y. Emelyanov1 and S. G. Gorokhova2
11 Sobolev Institute of Mathematics, Novosibirsk, Russia
22 Southern Mathematical Institute VSC RAS, Vladikavkaz, Russia
Keywords: ordered vector space, topological vector space, ordered Banach space, ordered Fréchet space, order-to-topology bounded operator
MSC2020: 46A40, 46B42, 47B02, 47B07, 47B60, 47B65
1 Introduction and preliminaries
Automatic boundedness of operators (enjoying some additional properties) plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g., [8, 9]). Another classical case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 7]. Apparently, the importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of such operators. The present paper is devoted to the question how far one can go with several other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 14, 15, 13, 16, 20, 18, 19, 22, 23, 25]).
Throughout, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$. We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS). Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$, and $a,b]=\{x\in X:a\leq x\leq b\}$ for the order interval in an OVS $X$, where $a,b\in X$. Throughout, the symbol ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$, and ${\rm L}(X,Y)$ for the space of bounded operators between NSs $X$ and $Y$, $B_{X}(a,r)=\{x\in X:d(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,d)$, and $B_{X}$ for $B_{X}(0,1)$ in NS $X$. It is proved in [15, Theorem 2.1] that order-to-norm bounded families of operators from a Banach lattice to an NS are equi-continuous. In [16, Theorem 2.1], it was extended to the OBS setting.
Let $X$ be a Banach space and $(Y,\tau)$ be a TVS. In absence of an additional structure in $X$, one can not say much about conditions providing boundedness of an operator $T:X\to Y$. Using of finite sets gives no proper subclass, because $T$ always carries finite sets into bounded ones. Whereas, the compactness returns us to the class of bounded operators, namely $T$ takes compact subsets into bounded sets iff $T$ is bounded. To see this, assume that $T$ is not bounded. Then, there exist an absorbing neighborhood $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ with $Tx_{n}\notin n^{2}U$ for all $n$. So, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ is not bounded.
The case of a Hilbert space in the domain was investigated in papers [18, 19], where it is proved in [19, Theorem 2.1] (see also [18, Lemma 1.2]) that
  1. -
[…]
Let us drop requirements of additional structures (like lattice operations, order, or inner products) on a space in the domain and consider:
Question 1.1.
Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it is bounded along normalized (unconditional) basic sequences of $X$?
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to Gowers and Maurey, who constructed a reflexive Banach space with no unconditional basic sequence [21]. For such a Banach space $X$, every linear functional on $X$ is bounded along bounded unconditional basic sequences simply because of absence of these sequences. So, take any unbounded functional $f:X\to\mathbb{R}$. Therefore, it makes sense to deal with the following:
Question 1.2.
How to describe pairs of Banach spaces $X$ and $Y$ such that every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
Although Question 1.2 is interesting and challenging, we are not going to study it in this paper.
Question 1.3.
How to describe the class of Banach lattices such that every linear functional on a Banach lattice bounded along bounded disjoint sequences is bounded?
The example constructed by Storozhuk in [24] suggests that many Banach lattices (e.g., discrete Banach function spaces) do not belong this class.
The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain. We mainly focus on Banach spaces in the domain and, when suitable, Fréchet spaces. The case of Fréchet spaces in the domain occurs recently for compactly classified operators along disjoint bounded sequences in [19].
The present paper is organized as follows. Section 2 is devoted to order-to-topology continuous and order-to-topology bounded operators between OBSs (or OVSs) and TVSs. Several conditions providing topological boundedness of such operators are investigated (as well as the study undertaking of the Uniform Boundedness Principle for collectively qualified families of them). Section 3 is devoted to conditions which provide boundedness of Levi and Lebesgue operators. For example, Theorem 3.2 tells that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded, and Theorem 3.11 states that if an OBS in the domain has a closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded.
For further unexplained terminology and notation, we refer to [1, 3, 4].
2 On boundedness of some operators between ordered TVSs
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by many authors (see, for example, [5, 15, 16, 18, 19, 22, 23]). Some of their results were extended to operators from an OBS to a TVS in [14]. It is important to find further conditions on operators in ordered TVSs which provide automatic boundedness (see, e.g. [3, 4, 7, 13, 14]).
In this section, we continue the research initiated in [14, 15, 13, 16]. Propositions 2.3, 2.4, 2.5, and 2.6 supply with several conditions providing topological boundedness of order-to-topology continuous and order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
Then, we adopt the definition of $\tau$-Levi and $\tau$-Lebesgue operators from [5, 20, 14] to the OVS-setting in the domain and investigate $\tau$-Lebesgue operators from an OBS to an NS. Lemma 3.10 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone. Theorem 3.11 states that if the domain has closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded. Theorem 3.3 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
This section is also aimed to introduce several notions and notation that will be used below.
Recall, that a net $(x_{\alpha})$ in an OVS $X$
  1. -
    order converges to $x$ (o-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
  2. -
    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
  3. -
    is order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
  4. -
    is eventually order bounded if there exist $[a,b]$ in $X$ and $\alpha_{0}$ such that $x_{\alpha}\in[a,b]$ for all $\alpha\geq\alpha_{0}$.
  5. -
    is o-Cauchy (resp., ru-Cauchy)) if the double net $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})$ o-converges (resp., ru-converges).
It is straightforward to see that o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) nets are eventually order bounded. In particular, o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) sequences are order bounded.
Definition 2.1.
Let $X$ and $Y$ be OVSs. An operator $T\in{\cal L}(X,Y)$ is
  1. $a)$
    order bounded if $T[a,b]$ is order bounded for every order interval $[a,b]$ in $X$ ((the set of such operators is denoted by ${\cal L}_{ob}(X,Y))$.
  2. $b)$
    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly $T\in{\cal L}_{oc}(X,Y))$.
  3. $c)$
    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ ((shortly $T\in{\cal L}_{rc}(X,Y))$.
Let $X$ be an OVS and $(Y,\tau)$ be a TVS. An operator $T\in{\cal L}(X,Y)$ is
  1. $d)$
    order-to-topology continuous ((or $T\in{\cal L}_{o{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{onc}(X,Y)$ when $Y$ is an NS)).
  2. $e)$
    ru-to-topology continuous ((or $T\in{\cal L}_{r{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{rnc}(X,Y)$ when $Y$ is an NS)).
  3. $f)$
    order-to-topology bounded ((or $T\in{\cal L}_{o{\tau}b}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal L}_{onb}(X,Y)$ when $Y$ is an NS)).
Clearly, the sets ${\cal L}_{ob}(X,Y))$, ${\cal L}_{rc}(X,Y))$, ${\cal L}_{onc}(X,Y)$, etc., given by Definition 2.1 are vector spaces.
[…]
Definition 2.2.
Let $X$ and $Y$ be OVSs. A subset $\cal T$} of ${\cal L}(X,Y)$ is
  1. $a)$
    collectively order bounded ((or ${\cal T}\in\text{\bf L}_{ob}(X,Y)$)) whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
Let $X$ be an OVS and $(Y,\tau)$ be a TVS. A subset $\cal T$} of ${\cal L}(X,Y)$ is
  1. $b)$
    collectively order-to-topology continuous $(or{\cal T}\in\text{\bf L}_{o{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
  2. $c)$
    collectively ru-to-topology continuous $(or{\cal T}\in\text{\bf L}_{r{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$ for an NS $Y)$.
  3. $d)$
    collectively order-to-topology bounded $(or{\cal T}\in\text{\bf L}_{o{\tau}b}(X,Y)$)) if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ for an NS $Y)$.
A set $\cal T$} of operators from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
  1. $e)$
    collectively continuous $(or{\cal T}\in\text{\bf L}_{c}(X,Y)$)) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
  2. $f)$
    collectively bounded $(or{\cal T}\in\text{\bf L}_{b}(X,Y)$)) if ${\cal T}U$ is $\tau$-bounded whenever $U$ is $\xi$-bounded.
The $\sigma$-versions of Definition 2.1 b–e) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with the corresponding notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
In many cases, collectively order-to-topology bounded sets agree with collectively ru-to-topology continuous ones.
Proposition 2.3.
([14, Theorem 2.1]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{r{\tau}c}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS. If, additionally, $X_{+}-X_{+}=X$ then $\text{\bf L}_{o{\tau}b}(X,Y)=\text{\bf L}_{r{\tau}c}(X,Y)$.
For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous (and if, additionally, $X_{+}$ is generating, ${\cal L}_{o{\tau}b}(X,Y)={\cal L}_{r{\tau}c}(X,Y)$).
The following theorem is a topological version of [16, Theorem 2.1].
Proposition 2.4.
([14, Theorem 2.3]) We have $\text{\bf L}_{o{\tau}c}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$, whenever $X$ is an Archimedean OVS, $X_{+}-X_{+}=X$, and $Y$ a TVS.
In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded.
The following proposition is a partial extension of [15, Theorem 2.1] and [16, Theorem 2.8].
Proposition 2.5.
([14, Theorem 2.5]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
It follows that every collectively order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [14], we collect few more consequences of Proposition 2.5 in the next proposition.
Proposition 2.6.
Let $X$ be an OBS with a closed generating cone.
  1. $i)$
    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets ([14, Corollary 2.9]).
  2. ii)ii)
    If the cone of $X$ is also normal and $Y$ is an NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ ([14, Corollary 2.12]).
  3. $iii)$
    If the cone of $X$ is also normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ ([14, Corollary 2.14]).
  4. $iv)$
    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded iff $Y_{+}$ is normal ([14, Proposition 2.15]).
Let $X$ be an OBS with a normal cone and $Y=(Y,\tau)$ be a Banach space with the weak topology.
  1. $v)$
    ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ ([14, Proposition 2.16]).
  2. $vi)$
    If the cone of $X$ is also closed and generating, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is bounded ([14, Corollary 2.18]).
The following theorem generalizes Proposition 2.5 to the setting of ordered Fréchet spaces. The idea of its proof below is motivated by the proof of [14, Theorem 2.5]. Recall that a Fréchet space $({\cal F},\tau)$ is an LCS, whose locally convex topology $\tau$ is metrizable by a complete metric.
Theorem 2.7.
We have $\text{\bf L}_{o{\tau}b}({\cal F},Y)\subseteq\text{\bf L}_{b}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a TVS.
Proof.
[…]
Corollary 2.8.
Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is topologically bounded.
Since $\text{\bf L}_{b}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$ whenever $X$ is an ordered TVS with a normal cone and $Y$ is a TVS, the following result follows from Proposition 2.3 and Theorem 2.7.
Corollary 2.9.
[…]
  1. $i)$
    $\text{\bf L}_{r{\tau}c}({\cal F},Y)=\text{\bf L}_{r{\tau}b}({\cal F},Y)=\text{\bf L}_{b}({\cal F},Y)$.
  2. ii)ii)
    ${\cal L}_{r{\tau}c}({\cal F},Y)={\cal L}_{o{\tau}b}({\cal F},Y)={\cal L}_{b}({\cal F},Y)$.
By [16, Theorem 2.4], ${\cal L}_{rc}(X,Y)={\cal L}_{ob}(X,Y)$ whenever OVSs $X$ and $Y$ both have generating cones. Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{o{\tau}b}(X,Y)$ if $Y$ is an ordered TVS with a normal cone, we obtain the following consequence of Corollary 2.9 ii)ii) providing conditions for automatic topological boundedness of ru-continuous operators.
Corollary 2.10.
[…]
3 On boundedness of Levi and Lebesgue operators between ordered TVSs
We use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 13, 20, 22] in the vector lattice-setting).
Definition 3.1.
An operator $T$ from an ordered TVS $(X,\tau)$ to an OVS $Y$ is
  1. $a)$
    quasi ($\sigma$-)Levi if $T$ takes $\tau$-bounded increasing nets ((sequences)) of $X$ to o-Cauchy ones (briefly, $T\in{\cal L}_{\text{\rm qLevi}}(X,Y)$ (resp., $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}(X,Y))$.
An operator $T$ from an OVS $X$ to an LCS $(Y,\xi)$ is
  1. $b)$
    $(\sigma$-))Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1]. The set of such operators is denoted by ${\cal L}_{Leb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{Leb}(X,Y)$).
  2. $c)$
    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1]. The set of such operators is denoted by ${\cal L}_{wLeb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{wLeb}(X,Y)$).
  3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf. Definition 2.1d)d)).
We prove that, under a rather mild condition, quasi $\sigma$-Levi operators are topologically bounded. This result generalizes [13, Theorem 3.1].
Theorem 3.2.
Let $T$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ordered TVS $(Y,\xi)$ with a normal cone. Then $T$ is topologically bounded.
Proof.
[…]
Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ iff $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1]. The same result holds also in the following more general setting.
Theorem 3.3.
Let $T$ be a positive operator from an OVS $X$ to an ordered TVS $(Y,\tau)$ with a normal cone. Then
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{o{\tau}c}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{o{\tau}c}(X,Y).$
Proof.
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Since obviously ${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{o{\tau}c}(X,Y)$.
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{o{\tau}c}(X,Y)$. Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{o{\tau}c}(X,Y)$. ∎
Corollary 3.4.
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{onc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{onc})_{+}(X,Y)$.
By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue iff $T$ is w-Lebesgue. We extend this fact to OVS-setting as follows.
Theorem 3.5.
Let $T$ be a positive order continuous operator from an OVS $X$ to a Hausdorff LCS $(Y,\tau)$ ordered by a normal cone. Then
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{wLeb}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{wLeb}(X,Y).$
Proof.
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Suppose $T\in{\cal L}_{wLeb}(X,Y)$ and take $x_{\alpha}\downarrow 0$ in $X$. Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$. By [4, Lemma 2.28], we conclude $Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$. Thus, $T\in{\cal L}_{Leb}(X,Y)$. ∎
Corollary 3.6.
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{wLeb})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{wLeb})_{+}(X,Y)$.
Question 3.7.
Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
Till the end of the section, we restrict ourselves to operators with a normed space in the co-domain. The following lemma is a direct consequence of Definition 3.1.
Lemma 3.8.
Let $X$ be an OVS and $Y$ be an LCS. Then
${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y),\ \ {\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y),$
${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
${\cal L}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcap{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcup{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y).$
Lemma 3.9.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\rm L}(X,Y)$.
Proof.
The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\rm L}(X,Y)$ follows from [16, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
By [16, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded. Accordingly to [16, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous. Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
  1. (*)
    ${\cal L}_{ob}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Another case of automatic boundedness of operators concerns the order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [16, Corollary 2.2], we have:
  1. (**)
    ${\cal L}_{oc}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$. An inspection of [16, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [22] without specifying the class of (just linear or continuous linear) operators. Automatic boundedness of such operators was investigated in recent paper [23].
We are going to establish some automatic boundedness conditions for order-to-weak continuous operators on OVSs. More precisely, we shall prove Theorem 3.11 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded. We begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
Lemma 3.10.
Let $X$ be a normal OBS and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$. In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$.
Proof.
Let $T:X\to Y$ be $\sigma$-w-Lebesgue. If $T$ is not order-to-norm bounded then $T[0,u]$ is not bounded for some $u\in X_{+}$. Since $X$ is normal then $[0,u]$ is bounded, say $\sup\limits_{x\in[0,u]}\|x\|\leq M$. Pick a sequence $(u_{n})$ in $[0,u]$ with $\|Tu_{n}\|\geq n2^{n}$, and set $y_{n}:=\|\cdot\|$-$\sum\limits_{k=n}^{\infty}2^{-k}u_{k}$ for $n\in\mathbb{N}$. Then $y_{n}\downarrow\geq 0$. Let $0\leq y_{0}\leq y_{n}$ for all $n\in\mathbb{N}$. Since $0\leq y_{0}\leq 2^{1-n}u$ and $\|2^{1-n}u\|\leq M2^{1-n}\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. Thus, $y_{n}\downarrow 0$. It follows from $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$ that $(Ty_{n})$ is w-null, and hence is norm bounded, that is absurd because $\|Ty_{n+1}-Ty_{n}\|=\|T(2^{-n}u_{n})\|\geq n$ for all $n\in\mathbb{N}$. ∎
Combining Lemmas 3.8, 3.9 and 3.10, we obtain.
Theorem 3.11.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(E,Y)\subseteq{\rm L}(E,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(E,Y)\subseteq{\rm L}(E,Y)$.
Corollary 3.12.
(cf. [5, Proposition 3.2], [23, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
The next result is a straightforward consequence of Lemma 3.8 and Theorem 3.11.
Corollary 3.13.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\rm L}(X,Y)$.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1. c)]). From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1)].
Definition 3.14.
The topology in an ordered TVS $(X,\tau)$ (the norm in an ONS $X$) is
  1. $a)$
    order continuous if $x_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{\alpha}\|\to 0$) for every net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
  2. $b)$
    $(\sigma$-))order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
In certain cases o-convergence is equivalent to ru-convergence.
Lemma 3.15.
Let $X$ be an OBS with a closed normal cone.
  1. $i)$
    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
  2. ii)ii)
    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
Proof.
First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Suppose $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$. Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$, $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$. Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$. Since $\big\|\frac{1}{n}u\big\|\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. It follows $\frac{1}{n}u\downarrow 0$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$.
$i)$ Let $X$ has order continuous norm and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Find an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$, and set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$.
ii)ii) The proof is similar. ∎
Theorem 3.16.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS.
  1. $i)$
    If $X$ has $\sigma$-order continuous norm then L^σ_Leb(X,Y)=L^σ_onc(X,Y) and L^σ_wLeb(X,Y)=L^σ_owc(X,Y).
  2. ii)ii)
    If $X$ has order continuous norm then L_Leb(X,Y)=L_onc(X,Y) and L_wLeb(X,Y)=L_owc(X,Y).
Proof.
We prove only $i)$ as the proof of ii)ii) is similar. In view of Lemma 3.8, we need to prove ${\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$ and ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y)$. Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. Then $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$, by Lemma 3.15, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\rm L}(X,Y)$ by Theorem 3.11, and hence $\|Tx_{n}\|\to 0$. If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$. Then again $T\in{\rm L}(X,Y)$ by Theorem 3.11. So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$. ∎
The next corollary is a partial extension of results of [22, Lemma 1], [5, Lemma 2.1], [25, Theorem 3], and [23, Proposition 1.5] to the OVS setting.
Corollary 3.17.
Let $E$ be a Banach lattice with order continuous norm and $Y$ be an NS. Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and
${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
Finally, we give some conditions on order bounded operators to be order-to-norm continuous.
Theorem 3.18.
Let $X$ be an OBS with a closed normal cone and $Y$ be an ONS with a generating normal cone.
  1. $i)$
    If $X$ has $\sigma$-order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$.
  2. ii)ii)
    If $X$ has order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{onc}(X,Y)$.
Proof.
We prove only $i)$ as the proof of ii)ii) is similar. Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.15, $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$. It follows from [16, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{(ru)}}0$. The normality of the cone $Y_{+}$ provides $\|Tx_{n}\|\to 0$. ∎

### 16 Apr 2026 v3

@@
-On automatic boundedness of operators between ordered and topological vector spaces
+On automatic boundedness of operators from ordered to topological vector spaces
 Abstract
-We study topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces.
-Levi and Lebesgue operators are investigated.
+We study continuity and topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces.
+The Levi operators and Lebesgue operators are investigated.
 E. Y. Emelyanov1 and S. G. Gorokhova2
 11 Sobolev Institute of Mathematics, Novosibirsk, Russia
-22 Southern Mathematical Institute VSC RAS, Vladikavkaz, Russia
-Keywords: ordered vector space, topological vector space, ordered Banach space, ordered Fréchet space, order-to-topology bounded operator
+22 Southern Mathematical Institute of the Vladikavkaz Scientific Center of the Russian Academy of Sciences, Vladikavkaz, Russia
+Keywords: ordered vector space, topological vector space, Banach space, Fréchet space, order-to-topology bounded operator
 MSC2020: 46A40, 46B42, 47B02, 47B07, 47B60, 47B65
 1 Introduction and preliminaries
-Automatic boundedness of operators (enjoying some additional properties) plays significant role in analysis.
-For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g., [8, 9]).
-Another classical case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 7].
-Apparently, the importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of such operators.
-The present paper is devoted to the question how far one can go with several other kinds of operators from a Banach space to a topological vector space.
+Automatic boundedness of operators enjoying some additional properties plays significant role in analysis.
+For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g.
+[7, 8]).
+Another case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 6].
+Apparently, importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of these operators.
+The present paper is devoted to the question how far one can go with other kinds of operators from a Banach space to a topological vector space.
 This question has attracted an attention in several recent papers (see, for example, [5, 11, 12, 13, 15, 14, 17, 18, 21, 22, 24]).
-Throughout, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$.
-We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS).
-Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$, and $a,b]=\{x\in X:a\leq x\leq b\}$ for the order interval in an OVS $X$, where $a,b\in X$.
-Throughout, the symbol ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$, and ${\rm L}(X,Y)$ for the space of bounded operators between NSs $X$ and $Y$, $B_{X}(a,r)=\{x\in X:d(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,d)$, and $B_{X}$ for $B_{X}(0,1)$ in NS $X$.
-It is proved in [15, Theorem 2.1] that order-to-norm bounded families of operators from a Banach lattice to an NS are equi-continuous.
-In [16, Theorem 2.1], it was extended to the OBS setting.
-Let $X$ be a Banach space and $(Y,\tau)$ be a TVS.
-In absence of an additional structure in $X$, one can not say much about conditions providing boundedness of an operator $T:X\to Y$.
-Using of finite sets gives no proper subclass, because $T$ always carries finite sets into bounded ones.
-Whereas, the compactness returns us to the class of bounded operators, namely $T$ takes compact subsets into bounded sets iff $T$ is bounded.
-To see this, assume that $T$ is not bounded.
-Then, there exist an absorbing neighborhood $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ with $Tx_{n}\notin n^{2}U$ for all $n$.
-So, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ is not bounded.
-The case of a Hilbert space in the domain was investigated in papers [18, 19], where it is proved in [19, Theorem 2.1] (see also [18, Lemma 1.2]) that
+Throughout, vector spaces are real, and operators are linear.
+We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively).
+Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS.
+It is proved in [13, Theorem 2.1] that each order-to-norm bounded family of operators from a Banach lattice to an NS is equi-continuous.
+In [15, Theorem 2.1], this result is extended to the OBS setting.
+Let $T:X\to Y$ be an operator from an NS $X$ to a TVS $(Y,\tau)$.
+In absence of additional structures in $X$, one cannot say much about conditions providing topological boundedness/continuity of $T$.
+For example, using of finite sets gives no proper subclass, because $T$ always carries finite sets into $\tau$-bounded sets.
+On the other hand, usage of compactness returns us to bounded operators.
+Indeed, if $T$ is $\tau$-bounded then $T$ takes compact sets into bounded sets.
+On the other hand, if $T$ is not $\tau$-bounded then there exist a circled $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ such that $Tx_{n}\notin n^{2}U$ for all $n\in\mathbb{N}$.
+In this case, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ under $T$ is not $\tau$-bounded.
+The case of a Hilbert space in the domain was investigated in papers [17, 18], where it is proved in [18, Theorem 2.1] (see also [17, Lemma 1.2]) that:
   1. -
 […]
-Let us drop requirements of additional structures (like lattice operations, order, or inner products) on a space in the domain and consider:
+Let us drop requirements of additional structures (e.g., lattice operations, order, or inner products) on a space in the domain and consider:
 Question 1.1.
 Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it is bounded along normalized (unconditional) basic sequences of $X$?
-The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to Gowers and Maurey, who constructed a reflexive Banach space with no unconditional basic sequence [21].
-For such a Banach space $X$, every linear functional on $X$ is bounded along bounded unconditional basic sequences simply because of absence of these sequences.
-So, take any unbounded functional $f:X\to\mathbb{R}$.
-Therefore, it makes sense to deal with the following:
+The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded along bounded unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space).
+Therefore, it makes sense to deal with the following questions.
 Question 1.2.
-How to describe pairs of Banach spaces $X$ and $Y$ such that every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
-Although Question 1.2 is interesting and challenging, we are not going to study it in this paper.
+Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
 Question 1.3.
 How to describe the class of Banach lattices such that every linear functional on a Banach lattice bounded along bounded disjoint sequences is bounded?
-The example constructed by Storozhuk in [24] suggests that many Banach lattices (e.g., discrete Banach function spaces) do not belong this class.
+An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to this class.
 The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain.
-We mainly focus on Banach spaces in the domain and, when suitable, Fréchet spaces.
-The case of Fréchet spaces in the domain occurs recently for compactly classified operators along disjoint bounded sequences in [19].
-The present paper is organized as follows.
-Section 2 is devoted to order-to-topology continuous and order-to-topology bounded operators between OBSs (or OVSs) and TVSs.
-Several conditions providing topological boundedness of such operators are investigated (as well as the study undertaking of the Uniform Boundedness Principle for collectively qualified families of them).
-Section 3 is devoted to conditions which provide boundedness of Levi and Lebesgue operators.
-For example, Theorem 3.2 tells that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded, and Theorem 3.11 states that if an OBS in the domain has a closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded.
-For further unexplained terminology and notation, we refer to [1, 3, 4].
+We mainly focus on Banach spaces in the domain and (when suitable) on Fréchet spaces.
+The case of Fréchet spaces in the domain occurred recently in study of compactly classified operators along disjoint bounded sequences [18].
+This paper is organized as follows.
+In Section 2, we study order-to-topology continuous/bounded operators between OBSs (or, also, OVSs) and TVSs.
+Conditions providing continuity of such operators are investigated.
+Section 3 is devoted to conditions providing boundedness/continuity of the Levi operators and the Lebesgue operators.
+For instance, Theorem 3.3 tells us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded.
+Theorem 3.12 states that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded.
+For further unexplained terminology about TVSs and OVSs, we refer the reader to [1, 3, 4].
 2 On boundedness of some operators between ordered TVSs
-Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by many authors (see, for example, [5, 15, 16, 18, 19, 22, 23]).
+Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by several authors (see, e.g., [5, 13, 15, 17, 18, 21, 22]).
 Some of their results were extended to operators from an OBS to a TVS in [12].
-It is important to find further conditions on operators in ordered TVSs which provide automatic boundedness (see, e.g.
-[3, 4, 7, 13, 14]).
+It is important to find further conditions on operators in ordered TVSs which provide automatic continuity/boundedness.
 In this section, we continue the research initiated in [12, 13, 11, 15].
-Propositions 2.3, 2.4, 2.5, and 2.6 supply with several conditions providing topological boundedness of order-to-topology continuous and order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
-Then, we adopt the definition of $\tau$-Levi and $\tau$-Lebesgue operators from [5, 20, 14] to the OVS-setting in the domain and investigate $\tau$-Lebesgue operators from an OBS to an NS.
+Propositions 2.3, 2.4, and 2.5, provides several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
+Then, we study $\tau$-Levi and $\tau$-Lebesgue operators following [5, 19, 12] in more general OVS-setting in the domain.
 Lemma 3.11 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone.
-Theorem 3.11 states that if the domain has closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded.
+Theorem 3.12 states that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded.
 Theorem 3.4 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
-This section is also aimed to introduce several notions and notation that will be used below.
-Recall, that a net $(x_{\alpha})$ in an OVS $X$
+We need several further notions and notation. Recall that a net $(x_{\alpha})$ in an OVS $X$
   1. -
@@
   2. -
-    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
+    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
   3. -
@@
   4. -
-    is eventually order bounded if there exist $[a,b]$ in $X$ and $\alpha_{0}$ such that $x_{\alpha}\in[a,b]$ for all $\alpha\geq\alpha_{0}$.
+    is eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
   5. -
-    is o-Cauchy (resp., ru-Cauchy)) if the double net $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})$ o-converges (resp., ru-converges).
-It is straightforward to see that o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) nets are eventually order bounded.
-In particular, o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) sequences are order bounded.
+    is o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{ru}}0$).
+It is straightforward to see that every o-convergent and every ru-convergent (hence, o-Cauchy and ru-Cauchy) net is eventually order bounded.
+In particular, o-convergent and ru-convergent (hence, also o-Cauchy and ru-Cauchy) sequences are order bounded.
 Definition 2.1.
-Let $X$ and $Y$ be OVSs. An operator $T\in{\cal L}(X,Y)$ is
+An operator $T$ from an OVS $X$ to an OVS $Y$ is
   1. $a)$
-    order bounded if $T[a,b]$ is order bounded for every order interval $[a,b]$ in $X$ ((the set of such operators is denoted by ${\cal L}_{ob}(X,Y))$.
+    order bounded if $T[a,b]$ is order bounded for every $a,b]\subseteq X$ ((shortly, $T\in{\cal L}_{ob}(X,Y))$.
   2. $b)$
-    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly $T\in{\cal L}_{oc}(X,Y))$.
+    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly, $T\in{\cal L}_{oc}(X,Y))$.
   3. $c)$
-    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ ((shortly $T\in{\cal L}_{rc}(X,Y))$.
-Let $X$ be an OVS and $(Y,\tau)$ be a TVS. An operator $T\in{\cal L}(X,Y)$ is
+    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ ((shortly, $T\in{\cal L}_{rc}(X,Y))$.
+An operator $T$ from an OVS $X$ to a TVS $(Y,\tau)$ is
   1. $d)$
-    order-to-topology continuous ((or $T\in{\cal L}_{o{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{onc}(X,Y)$ when $Y$ is an NS)).
+    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{onc}(X,Y)$, when $Y$ is an NS)).
   2. $e)$
-    ru-to-topology continuous ((or $T\in{\cal L}_{r{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{rnc}(X,Y)$ when $Y$ is an NS)).
+    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{rnc}(X,Y)$, when $Y$ is an NS)).
   3. $f)$
-    order-to-topology bounded ((or $T\in{\cal L}_{o{\tau}b}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal L}_{onb}(X,Y)$ when $Y$ is an NS)).
-Clearly, the sets ${\cal L}_{ob}(X,Y))$, ${\cal L}_{rc}(X,Y))$, ${\cal L}_{onc}(X,Y)$, etc., given by Definition 2.1 are vector spaces.
-[…]
+    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write $T\in{\cal L}_{onb}(X,Y)$, when $Y$ is an NS)).
+An operator $T$ from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
+  1. $g)$
+    continuous $(T\in{\cal L}_{c}(X,Y))$ if $Tx_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
+  2. $e)$
+    topologically-bounded $(T\in{\cal L}_{tb}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$ ((we write ${\cal L}_{b}(X,Y)$, when $Y$ and $Y$ are NSs)).
+Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{onc}(X,Y)$, etc. given by Definition 2.1 are vector spaces.
+We recall several notions on collectively qualified sets of operators (cf.
+[9, 10, 12, 13, 17, 18, 15]).
+A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}\operatorname{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
 Definition 2.2.
-Let $X$ and $Y$ be OVSs. A subset $\cal T$} of ${\cal L}(X,Y)$ is
+A set $\cal T$} of operators from an OVS $X$ to an OVS $Y$ is
   1. $a)$
-    collectively order bounded ((or ${\cal T}\in\text{\bf L}_{ob}(X,Y)$)) whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
-Let $X$ be an OVS and $(Y,\tau)$ be a TVS. A subset $\cal T$} of ${\cal L}(X,Y)$ is
+    equi-order bounded $({\cal T}\in\text{\bf L}_{ob}(X,Y))$, whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
+A set $\cal T$} of operators from an OVS $X$ to a TVS $(Y,\tau)$ is
   1. $b)$
-    collectively order-to-topology continuous $(or{\cal T}\in\text{\bf L}_{o{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
+    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
   2. $c)$
-    collectively ru-to-topology continuous $(or{\cal T}\in\text{\bf L}_{r{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$ for an NS $Y)$.
+    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$, when $Y$ is an NS)).
   3. $d)$
-    collectively order-to-topology bounded $(or{\cal T}\in\text{\bf L}_{o{\tau}b}(X,Y)$)) if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ for an NS $Y)$.
+    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$, when $Y$ is an NS)).
 A set $\cal T$} of operators from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
   1. $e)$
-    collectively continuous $(or{\cal T}\in\text{\bf L}_{c}(X,Y)$)) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
+    equi-continuous $({\cal T}\in\text{\bf L}_{c}(X,Y))$ if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
   2. $f)$
-    collectively bounded $(or{\cal T}\in\text{\bf L}_{b}(X,Y)$)) if ${\cal T}U$ is $\tau$-bounded whenever $U$ is $\xi$-bounded.
-The $\sigma$-versions of Definition 2.1 b–e) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with the corresponding notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
-In many cases, collectively order-to-topology bounded sets agree with collectively ru-to-topology continuous ones.
+    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{tb}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded ((we write ${\cal T}\in\text{\bf L}_{b}(X,Y)$, when $X$ and $Y$ are NSs)).
+The sequential versions of Definition 2.1 b)–e), g) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
+In many cases, equi-order-to-topology bounded sets agree with equi-ru-to-topology continuous ones.
 Proposition 2.3.
-([14, Theorem 2.1]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{r{\tau}c}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS.
-If, additionally, $X_{+}-X_{+}=X$ then $\text{\bf L}_{o{\tau}b}(X,Y)=\text{\bf L}_{r{\tau}c}(X,Y)$.
-For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous (and if, additionally, $X_{+}$ is generating, ${\cal L}_{o{\tau}b}(X,Y)={\cal L}_{r{\tau}c}(X,Y)$).
-The following theorem is a topological version of [16, Theorem 2.1].
+([12, Theorem 2.1]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS.
+If additionally $X_{+}$ is generating then $\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{rtc}(X,Y)$.
+For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$.
+The following proposition is a topological counterpart of [15, Theorem 2.1].
 Proposition 2.4.
-([14, Theorem 2.3]) We have $\text{\bf L}_{o{\tau}c}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$, whenever $X$ is an Archimedean OVS, $X_{+}-X_{+}=X$, and $Y$ a TVS.
+([12, Theorem 2.3]) $\text{\bf L}_{otc}(X,Y)\subseteq\text{\bf L}_{otb}(X,Y)$, whenever $X$ is an Archimedean OVS with a generating cone, and $Y$ is a TVS.
 In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded.
-The following proposition is a partial extension of [15, Theorem 2.1] and [16, Theorem 2.8].
+The following proposition is a partial extension of [13, Theorem 2.1] and [15, Theorem 2.8].
 Proposition 2.5.
-([14, Theorem 2.5]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
-It follows that every collectively order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded.
+([12, Theorem 2.5]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
+It follows that every equi-order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded.
 In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous.
-Following [14], we collect few more consequences of Proposition 2.5 in the next proposition.
-Proposition 2.6.
-Let $X$ be an OBS with a closed generating cone.
+Following [12], we also collect few more statements below.
+Let $X$ be an ordered Banach space with a closed generating cone.
   1. $i)$
-    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets ([14, Corollary 2.9]).
+    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets [12, Corollary 2.9].
   2. ii)ii)
-    If the cone of $X$ is also normal and $Y$ is an NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ ([14, Corollary 2.12]).
+    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
   3. $iii)$
-    If the cone of $X$ is also normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ ([14, Corollary 2.14]).
+    If $X_{+}$ is normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ [12, Corollary 2.14].
   4. $iv)$
-    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded iff $Y_{+}$ is normal ([14, Proposition 2.15]).
-Let $X$ be an OBS with a normal cone and $Y=(Y,\tau)$ be a Banach space with the weak topology.
+    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, where $X^{\sim}$ is an order dual of $X$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded if and only if $Y_{+}$ is normal [12, Proposition 2.15].
   5. $v)$
-    ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ ([14, Proposition 2.16]).
+    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ [12, Proposition 2.16].
   6. $vi)$
-    If the cone of $X$ is also closed and generating, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is bounded ([14, Corollary 2.18]).
-The following theorem generalizes Proposition 2.5 to the setting of ordered Fréchet spaces.
-The idea of its proof below is motivated by the proof of [14, Theorem 2.5].
-Recall that a Fréchet space $({\cal F},\tau)$ is an LCS, whose locally convex topology $\tau$ is metrizable by a complete metric.
+    If $X_{+}$ is closed generating and normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is topologically bounded [12, Corollary 2.18].
+[…]
 Theorem 2.6.
-We have $\text{\bf L}_{o{\tau}b}({\cal F},Y)\subseteq\text{\bf L}_{b}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a TVS.
+$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{c}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
 Proof.
+Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{c}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
+$V_{n}:=B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}-B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}.$
+Clearly, $\{B_{\cal F}(0,2^{-2n}):n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of $\tau$-closed sets and $B_{\cal F}(0,2^{-2(n+1)})+B_{\cal F}(0,2^{-2(n+1)})\subseteq B_{\cal F}(0,2^{-2n})$ holds for each $n$.
+Furthermore, each $B_{\cal F}(0,2^{-2n})$ is circled.
+To see this, it suffices to show that $p(x,0)\leq 2^{-2n}$ implies $p(tx,0)\leq 2^{-2n}$ whenever $|t|\leq 1$.
+So, let $p(x,0)\leq 2^{-2n}$ and $|t|\leq 1$.
+Since $\frac{|t|a}{1+|t|a}\leq\frac{a}{1+a}$ for every $a\in\mathbb{R}_{+}$, we obtain $p(tx,0)=\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(tx)}{1+p_{k}(tx)}=\sum\limits_{k=1}^{\infty}\frac{2^{-k}|t|p_{k}(x)}{1+|t|p_{k}(x)}\leq\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(x)}{1+p_{k}(x)}=p(x,0)\leq 2^{-2n}$.
+By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of (not necessarily open) circled subsets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
+[…]
 […]
 Corollary 2.7.
-Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is topologically bounded.
-Since $\text{\bf L}_{b}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$ whenever $X$ is an ordered TVS with a normal cone and $Y$ is a TVS, the following result follows from Proposition 2.3 and Theorem 2.7.
+Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a topological vector space is continuous.
+Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$ if $Y$ is a TVS with a normal cone [4, Lemma 2.22], we obtain the following conditions for automatic continuity of order bounded operators.
 Corollary 2.8.
-[…]
-  1. $i)$
-    $\text{\bf L}_{r{\tau}c}({\cal F},Y)=\text{\bf L}_{r{\tau}b}({\cal F},Y)=\text{\bf L}_{b}({\cal F},Y)$.
-  2. ii)ii)
-    ${\cal L}_{r{\tau}c}({\cal F},Y)={\cal L}_{o{\tau}b}({\cal F},Y)={\cal L}_{b}({\cal F},Y)$.
-By [16, Theorem 2.4], ${\cal L}_{rc}(X,Y)={\cal L}_{ob}(X,Y)$ whenever OVSs $X$ and $Y$ both have generating cones.
-Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{o{\tau}b}(X,Y)$ if $Y$ is an ordered TVS with a normal cone, we obtain the following consequence of Corollary 2.9 ii)ii) providing conditions for automatic topological boundedness of ru-continuous operators.
-Corollary 2.10.
-[…]
-3 On boundedness of Levi and Lebesgue operators between ordered TVSs
-We use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 13, 20, 22] in the vector lattice-setting).
+Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered locally convex space with a normal cone is continuous.
+3 On continuity of the Levi and the Lebesgue operators
+We shall use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see also [5, 11, 19, 21] in the vector lattice-setting).
 Definition 3.1.
@@
     $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf.
-Definition 2.1d)d)).
-We prove that, under a rather mild condition, quasi $\sigma$-Levi operators are topologically bounded. This result generalizes [13, Theorem 3.1].
+The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf., Definition 2.1d)d)).
+We prove that, under rather mild conditions, quasi $\sigma$-Levi operators are continuous. This result generalizes [11, Theorem 3.1].
 Theorem 3.2.
-Let $T$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ordered TVS $(Y,\xi)$ with a normal cone.
-Then $T$ is topologically bounded.
+Every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
 Proof.
+Let $T:E\to F$ be a quasi $\sigma$-Levi operator from a BL $E$ to an NL $F$.
+On the way to a contradiction, suppose $T$ is not continuous.
+Then, there exists a sequence $(x_{n})$ in $B_{X}$ such that $\|Tx_{n}\|\to\infty$.
+Without loss of generality, suppose $(x_{n})\subset(B_{X})^{+}$ and $\|Tx_{n}\|\geq n2^{n}$ for every $n$.
+Set $u=\sum\limits_{k=1}^{\infty}2^{-k}x_{k}$ and $u_{n}=\sum\limits_{k=1}^{n}2^{-k}x_{k}$ for each $n\in\mathbb{N}$.
+Then $0\leq u_{n}\uparrow$ and $\|u_{n}\|\leq 1$ for all $n$.
+Since $T\in\text{\rm L}_{\text{\rm qLevi}}(E,F)$, $(Tu_{n})$ is an o-Cauchy sequence in $F$, and hence the sequence $(Tu_{n})$ is order bounded.
+Thus, $(Tu_{n})$ is also norm bounded.
+It is absurd because $\|Tu_{n+1}-Tu_{n}\|=\|T2^{-(n+1)}x_{n+1}\|\geq n+1$ for all $n$.
+∎
+Theorem 3.3.
+Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered normed space with a normal cone is continuous.
+Proof.
+Let $T:{\cal F}\to Y$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ONS $Y$ with a normal cone.
+Toward a contradiction, suppose that $T$ is not continuous.
 […]
-Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ iff $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1].
+Letting $x_{n}:=2^{n}z_{n}$, denote $u_{i}=\sum\limits_{k=1}^{i}x_{k}$ for each $i\in\mathbb{N}$.
+Clearly, $(u_{i})$ is an increasing sequence.
+Take an $l\in\mathbb{N}$.
+It is straightforward to see that, for each $m\in\mathbb{N}$,
+$\sum\limits_{k=l+1}^{m+l+1}x_{k}=\sum\limits_{k=l+1}^{m+l+1}2^{k}z_{k}\in\sum\limits_{k=l+1}^{m+l+1}2^{k}B_{\cal F}(0,2^{-2k})\subseteq\sum\limits_{k=0}^{m}2^{-k}B_{\cal F}(0,2^{-(l+1)})\subseteq B_{\cal F}(0,2^{-l}).$
+Find an $N>0$ satisfying $\sum\limits_{k=1}^{j}x_{k}\in NB_{\cal F}(0,2^{-l})$ for j=1,,lj=1,...,l. Then, for every $i\in\mathbb{N}$,
+$u_{i}=\sum\limits_{k=1}^{i}x_{k}=\sum\limits_{k=1}^{\min(i,l)}x_{k}+\sum\limits_{k=\min(i,l)+1}^{i}x_{k}\in(N+1)B_{\cal F}(0,2^{-l}).$
+As $\{B_{\cal F}(0,2^{-l})\}_{l\in\mathbb{N}}$ is a $\tau$-neighborhood base of zero, the sequence $(u_{i})$ is $\tau$-bounded.
+It follows from $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}({\cal F},Y)$ that $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded.
+Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is also norm bounded.
+It is absurd because
+$\|Tu_{i+1}-Tu_{i}\|=\|Tx_{i+1}\|=\|T(2^{i+1}z_{i+1})\|\geq 2^{i+1}M$
+for all $i\in\mathbb{N}$. ∎
+Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ if and only if $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1].
 The same result holds also in the following more general setting.
@@
 Let $T$ be a positive operator from an OVS $X$ to an ordered TVS $(Y,\tau)$ with a normal cone. Then
-$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{o{\tau}c}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{o{\tau}c}(X,Y).$
+$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{otc}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{otc}(X,Y).$
 Proof.
 As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence.
-Since obviously ${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{o{\tau}c}(X,Y)$.
-On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{o{\tau}c}(X,Y)$.
+Since obviously ${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{otc}(X,Y)$.
+On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{otc}(X,Y)$.
 Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$.
@@
 It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$.
-Thus, $T\in{\cal L}_{o{\tau}c}(X,Y)$.
+Thus, $T\in{\cal L}_{otc}(X,Y)$.
 ∎
@@
 Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{onc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{onc})_{+}(X,Y)$.
-By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue iff $T$ is w-Lebesgue. We extend this fact to OVS-setting as follows.
+By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue if and only if $T$ is w-Lebesgue.
+We extend this fact to the OVS-setting as follows.
 Theorem 3.6.
@@
 Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{wLeb})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{wLeb})_{+}(X,Y)$.
-Question 3.7.
-Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
-Till the end of the section, we restrict ourselves to operators with a normed space in the co-domain. The following lemma is a direct consequence of Definition 3.1.
+Question 3.8.
+Are the conclusions of Corollaries 3.5 and 3.7 hold true for arbitrary (not necessarily positive) operators?
+The following lemma is a direct consequence of Definition 3.1.
 Lemma 3.9.
@@
 ${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y),\ \ {\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y),$
-${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
-${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
+${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
+${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
 ${\cal L}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcap{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcup{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y).$
 Lemma 3.10.
-Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\rm L}(X,Y)$.
+Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\cal L}_{b}(X,Y)$.
 Proof.
-The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\rm L}(X,Y)$ follows from [16, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
-By [16, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded.
-Accordingly to [16, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous.
+The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
+By [15, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded.
+Accordingly to [15, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous.
 Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
   1. (*)
-    ${\cal L}_{ob}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
+    ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
 Another case of automatic boundedness of operators concerns the order continuity.
-Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [16, Corollary 2.2], we have:
+Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have:
   1. (**)
-    ${\cal L}_{oc}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
+    ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
 Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$.
-An inspection of [16, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
+An inspection of [15, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
 It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [21] without specifying the class of (just linear or continuous linear) operators.
@@
 ∎
-Combining Lemmas 3.8, 3.9 and 3.10, we obtain.
+Combining Lemmas 3.9, 3.10, and 3.11, we obtain.
 Theorem 3.12.
-Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(E,Y)\subseteq{\rm L}(E,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(E,Y)\subseteq{\rm L}(E,Y)$.
+Let $X$ be an ordered Banach space with a closed generating normal cone and let $Y$ be a normed space.
+Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
+In particular, ${\cal L}^{\sigma}_{ows}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
 Corollary 3.13.
-(cf. [5, Proposition 3.2], [23, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
+(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
 The next result is a straightforward consequence of Lemma 3.9 and Theorem 3.12.
 Corollary 3.14.
-Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS.
-Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\rm L}(X,Y)$.
+Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS.
+Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\cal L}_{b}(X,Y)$.
 Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1.
 c)]).
-From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1)].
+From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1].
 Definition 3.15.
@@
   2. $b)$
-    $(\sigma$-))order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+    $\sigma$-order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
 In certain cases o-convergence is equivalent to ru-convergence.
 Lemma 3.16.
-Let $X$ be an OBS with a closed normal cone.
+Let $X$ be an OBS with a closed normal cone. The following assertions hold.
   1. $i)$
-    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
+    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
   2. ii)ii)
-    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
+    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
 Proof.
-First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-Suppose $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$.
-Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$, $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
+First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+Suppose $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
+Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$ with $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
 Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$.
@@
 Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$.
-Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$.
+Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
 ii)ii) The proof is similar. ∎
 Theorem 3.17.
-Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS.
+Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. The following assertions hold.
   1. $i)$
@@
 Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-Then $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$, by Lemma 3.15, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
-If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\rm L}(X,Y)$ by Theorem 3.11, and hence $\|Tx_{n}\|\to 0$.
+Then $x_{n}\operatorname{\xrightarrow[]{ru}}0$, by Lemma 3.16, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
+If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12, and hence $\|Tx_{n}\|\to 0$.
 If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$.
-Then again $T\in{\rm L}(X,Y)$ by Theorem 3.11.
+Then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12.
 So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$.
 ∎
-The next corollary is a partial extension of results of [22, Lemma 1], [5, Lemma 2.1], [25, Theorem 3], and [23, Proposition 1.5] to the OVS setting.
+The next corollary is a partial extension of results of [21, Lemma 1], [5, Lemma 2.1], [24, Theorem 3], and [22, Proposition 1.5] to the OVS setting.
 Corollary 3.18.
-Let $E$ be a Banach lattice with order continuous norm and $Y$ be an NS.
-Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and
-${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
+Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space.
+Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
 Finally, we give some conditions on order bounded operators to be order-to-norm continuous.
 Theorem 3.19.
-Let $X$ be an OBS with a closed normal cone and $Y$ be an ONS with a generating normal cone.
+Let $X$ be an ordered Banach space with a closed normal cone and $Y$ an ordered normed space with a generating normal cone. The following assertions hold.
   1. $i)$
@@
 Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-By Lemma 3.15, $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$.
-It follows from [16, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{(ru)}}0$.
-The normality of the cone $Y_{+}$ provides $\|Tx_{n}\|\to 0$.
+By Lemma 3.16, $x_{n}\operatorname{\xrightarrow[]{ru}}0$.
+It follows from [15, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{ru}}0$.
+The normality of the cone $Y_{+}$ ensures $\|Tx_{n}\|\to 0$.
 ∎

### 20 Apr 2026 v4

@@
-On automatic boundedness of operators from ordered to topological vector spaces
+On automatic continuity of operators from ordered to topological vector spaces
 Abstract
-We study continuity and topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces.
-The Levi operators and Lebesgue operators are investigated.
+We study continuity and boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered to topological vector spaces.
+Several results on automatic continuity of operators from ordered Fréchet spaces to topological vector spaces are included.
+Levi and Lebesgue operators especially are investigated.
 E. Y. Emelyanov1 and S. G. Gorokhova2
@@
 1 Introduction and preliminaries
-Automatic boundedness of operators enjoying some additional properties plays significant role in analysis.
+Automatic continuity of operators enjoying some additional properties plays significant role in analysis.
 For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g.
@@
 Apparently, importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of these operators.
-The present paper is devoted to the question how far one can go with other kinds of operators from a Banach space to a topological vector space.
+This paper is devoted to the study of the question how far one can go with other kinds of operators from a Banach space to a topological vector space.
 This question has attracted an attention in several recent papers (see, for example, [5, 11, 12, 13, 15, 14, 17, 18, 21, 22, 24]).
 Throughout, vector spaces are real, and operators are linear.
-We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively).
-Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS.
+When suitable, we abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively).
+In what follows, symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval in an OVS $X$; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS.
 It is proved in [13, Theorem 2.1] that each order-to-norm bounded family of operators from a Banach lattice to an NS is equi-continuous.
 In [15, Theorem 2.1], this result is extended to the OBS setting.
-Let $T:X\to Y$ be an operator from an NS $X$ to a TVS $(Y,\tau)$.
+In [12], the further extension is obtained in the case of a TVS in the co-domain.
+In the present paper, we prove Theorem 2.6 that generalizes [12, Theorem 2.1] to operators from an ordered Fréchet space with a closed generating cone to a TVS.
+Let $T:X\to Y$ be an operator from a normed space $X$ to a TVS $(Y,\tau)$.
 In absence of additional structures in $X$, one cannot say much about conditions providing topological boundedness/continuity of $T$.
@@
 Question 1.1.
-Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it is bounded along normalized (unconditional) basic sequences of $X$?
-The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded along bounded unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space).
+Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it takes normalized (unconditional) basic sequences of $X$ onto bounded sequences of $Y$?
+The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded on normalized unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space).
 Therefore, it makes sense to deal with the following questions.
-Question 1.2.
-Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
-Question 1.3.
-How to describe the class of Banach lattices such that every linear functional on a Banach lattice bounded along bounded disjoint sequences is bounded?
-An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to this class.
-The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain.
-We mainly focus on Banach spaces in the domain and (when suitable) on Fréchet spaces.
+  1. -
+    Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ that is bounded on normalized unconditional basic sequences of $X$ is bounded?
+  2. -
+    How to describe the class $S$ of Banach lattices such that every linear functional on a Banach lattice $X$ from $S$ is bounded, whenever it is bounded on every normalized disjoint sequence of $X$?
+An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to the class $S$.
+The present paper deals with operators between vector spaces, where the order structure is assumed in the domain or/and codomain.
+We focus on Banach spaces and, when suitable, Fréchet spaces in the domain.
 The case of Fréchet spaces in the domain occurred recently in study of compactly classified operators along disjoint bounded sequences [18].
 This paper is organized as follows.
-In Section 2, we study order-to-topology continuous/bounded operators between OBSs (or, also, OVSs) and TVSs.
-Conditions providing continuity of such operators are investigated.
-Section 3 is devoted to conditions providing boundedness/continuity of the Levi operators and the Lebesgue operators.
-For instance, Theorem 3.3 tells us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded.
-Theorem 3.12 states that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded.
-For further unexplained terminology about TVSs and OVSs, we refer the reader to [1, 3, 4].
-2 On boundedness of some operators between ordered TVSs
+In Section 2, we survey briefly some recently published results on order-to-topology continuous/bounded operators between OBSs and TVSs.
+Then, we prove Theorem 2.6 on continuity of order-to-topology bounded operators from an ordered Fréchet space with a closed generating cone to a TVS.
+In Section 3, we study conditions providing continuity of Levi and Lebesgue operators.
+For instance, we prove: Theorem 3.2 telling us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is continuous; and Theorem 3.12 saying that each $\sigma$-w-Lebesgue operator from an ordered Banach space with a closed generating normal cone to a normed space is bounded.
+For further unexplained terminology and notation, we refer the reader to [1, 3, 4].
+2 On continuity of some operators between ordered topological vector spaces
 Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by several authors (see, e.g., [5, 13, 15, 17, 18, 21, 22]).
@@
 In this section, we continue the research initiated in [12, 13, 11, 15].
-Propositions 2.3, 2.4, and 2.5, provides several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
-Then, we study $\tau$-Levi and $\tau$-Lebesgue operators following [5, 19, 12] in more general OVS-setting in the domain.
-Lemma 3.11 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone.
-Theorem 3.12 states that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded.
-Theorem 3.4 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
-We need several further notions and notation. Recall that a net $(x_{\alpha})$ in an OVS $X$
-  1. -
-    order converges to $x$ (o-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
-  2. -
-    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
-  3. -
-    is order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
-  4. -
-    is eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
-  5. -
-    is o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{ru}}0$).
-It is straightforward to see that every o-convergent and every ru-convergent (hence, o-Cauchy and ru-Cauchy) net is eventually order bounded.
+First, we point out Propositions 2.4 and 2.5 providing several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
+Then, we prove Theorem 2.6 telling us that every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is continuous.
+We need several further notions and notation.
+Definition 2.1.
+A net $(x_{\alpha})$ in an ordered vector space $X$ is
+  1. $a)$
+    order convergent to $x$ (o-convergent to $x$, or $x_{\alpha}{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
+  2. $b)$
+    relative uniform convergent to $x$ (ru-convergent to $x$, or $x_{\alpha}{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
+  3. $c)$
+    order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
+  4. $d)$
+    eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
+  5. $e)$
+    o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}}){\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}}){\xrightarrow[]{ru}}0$).
+It is straightforward to see that every o-convergent and every ru-convergent (hence, every o-Cauchy and every ru-Cauchy) net is eventually order bounded.
 In particular, o-convergent and ru-convergent (hence, also o-Cauchy and ru-Cauchy) sequences are order bounded.
@@
   1. $a)$
-    order bounded if $T[a,b]$ is order bounded for every $a,b]\subseteq X$ ((shortly, $T\in{\cal L}_{ob}(X,Y))$.
+    order bounded $(T\in{\cal L}_{ob}(X,Y))$ if $T[a,b]$ is order bounded for every $a,b]\subseteq X$.
   2. $b)$
-    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly, $T\in{\cal L}_{oc}(X,Y))$.
+    order continuous $(T\in{\cal L}_{oc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}{\xrightarrow[]{o}}0$.
   3. $c)$
-    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ ((shortly, $T\in{\cal L}_{rc}(X,Y))$.
+    ru-continuous $(T\in{\cal L}_{rc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}{\xrightarrow[]{ru}}0$.
 An operator $T$ from an OVS $X$ to a TVS $(Y,\tau)$ is
   1. $d)$
-    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{onc}(X,Y)$, when $Y$ is an NS)).
+    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
   2. $e)$
-    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{rnc}(X,Y)$, when $Y$ is an NS)).
+    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
   3. $f)$
-    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write $T\in{\cal L}_{onb}(X,Y)$, when $Y$ is an NS)).
+    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$.
 An operator $T$ from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
   1. $g)$
-    continuous $(T\in{\cal L}_{c}(X,Y))$ if $Tx_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
+    continuous $(T\in{\cal L}_{tc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{\tau}}0$, whenever $x_{\alpha}{\xrightarrow[]{\xi}}0$.
   2. $e)$
-    topologically-bounded $(T\in{\cal L}_{tb}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$ ((we write ${\cal L}_{b}(X,Y)$, when $Y$ and $Y$ are NSs)).
-Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{onc}(X,Y)$, etc. given by Definition 2.1 are vector spaces.
+    topologically-bounded $(T\in{\cal L}_{b}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$.
+Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{otc}(X,Y)$, etc. given by Definition 2.2 are vector spaces.
 We recall several notions on collectively qualified sets of operators (cf.
 [9, 10, 12, 13, 17, 18, 15]).
-A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}\operatorname{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
+A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
 Definition 2.3.
@@
   1. $b)$
-    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
+    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$.
   2. $c)$
-    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$, when $Y$ is an NS)).
+    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$.
   3. $d)$
-    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$, when $Y$ is an NS)).
+    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$.
 A set $\cal T$} of operators from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
   1. $e)$
-    equi-continuous $({\cal T}\in\text{\bf L}_{c}(X,Y))$ if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
+    equi-continuous $({\cal T}\in\text{\bf L}_{tc}(X,Y))$ if ${\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}{\xrightarrow[]{\xi}}0$.
   2. $f)$
-    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{tb}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded ((we write ${\cal T}\in\text{\bf L}_{b}(X,Y)$, when $X$ and $Y$ are NSs)).
-The sequential versions of Definition 2.1 b)–e), g) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
-In many cases, equi-order-to-topology bounded sets agree with equi-ru-to-topology continuous ones.
+    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{b}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded.
+The sequential versions of Definition 2.2 b)–e), g) and Definition 2.3 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{tc}^{\sigma}(X,Y)$, etc.).
+We recall several recent results concerned the notions listed in Definitions 2.2 and 2.3.
 Proposition 2.4.
-([12, Theorem 2.1]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS.
+([12, Theorem 2.1]) If $X$ is an OVS and $Y$ is a TVS then $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$.
 If additionally $X_{+}$ is generating then $\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{rtc}(X,Y)$.
-For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$.
-The following proposition is a topological counterpart of [15, Theorem 2.1].
-Proposition 2.4.
-([12, Theorem 2.3]) $\text{\bf L}_{otc}(X,Y)\subseteq\text{\bf L}_{otb}(X,Y)$, whenever $X$ is an Archimedean OVS with a generating cone, and $Y$ is a TVS.
-In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded.
+In particular, each order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, then ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$.
 The following proposition is a partial extension of [13, Theorem 2.1] and [15, Theorem 2.8].
 Proposition 2.5.
-([12, Theorem 2.5]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
+([12, Theorem 2.5]) If $X$ is an OBS with a closed generating cone and $Y$ is a TVS then $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$.
 It follows that every equi-order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded.
 In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous.
-Following [12], we also collect few more statements below.
+Following [12], we collect few more statements below.
 Let $X$ be an ordered Banach space with a closed generating cone.
@@
   2. ii)ii)
-    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
+    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rtc}(X,Y)=\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
   3. $iii)$
@@
   5. $v)$
-    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ [12, Proposition 2.16].
+    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$ [12, Proposition 2.16].
   6. $vi)$
@@
 Theorem 2.6.
-$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{c}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
+$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{tc}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
 Proof.
-Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{c}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
+Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{tc}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
 $V_{n}:=B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}-B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}.$
@@
 Since $\frac{|t|a}{1+|t|a}\leq\frac{a}{1+a}$ for every $a\in\mathbb{R}_{+}$, we obtain $p(tx,0)=\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(tx)}{1+p_{k}(tx)}=\sum\limits_{k=1}^{\infty}\frac{2^{-k}|t|p_{k}(x)}{1+|t|p_{k}(x)}\leq\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(x)}{1+p_{k}(x)}=p(x,0)\leq 2^{-2n}$.
-By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of (not necessarily open) circled subsets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
+By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of circled sets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
 […]
@@
 Corollary 2.8.
-Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered locally convex space with a normal cone is continuous.
+Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
 3 On continuity of the Levi and the Lebesgue operators
-We shall use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see also [5, 11, 19, 21] in the vector lattice-setting).
+In this section, we study $\tau$-Levi and $\tau$-Lebesgue operators [5, 19, 12] in the more general OVS-setting.
+We prove Theorem 3.2 saying that a quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
+This theorem is a far going generalization of [11, Theorem 3.1], where the continuity is established for a quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice.
+We mention also: Lemma 3.11 saying that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has a normal cone; Theorem 3.12 saying that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded; and Theorem 3.3 providing conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
+We shall use below the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 11, 19, 21] in the vector lattice-setting).
 Definition 3.1.
@@
   1. $a)$
-    quasi ($\sigma$-)Levi if $T$ takes $\tau$-bounded increasing nets ((sequences)) of $X$ to o-Cauchy ones (briefly, $T\in{\cal L}_{\text{\rm qLevi}}(X,Y)$ (resp., $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}(X,Y))$.
+    quasi $\sigma$-Levi if $T$ takes $\tau$-bounded increasing sequences of $X$ to o-Cauchy ones.
 An operator $T$ from an OVS $X$ to an LCS $(Y,\xi)$ is
   1. $b)$
-    $(\sigma$-))Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1].
+    $(\sigma$-))Lebesgue if $Tx_{\alpha}{\xrightarrow[]{\tau}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$.
 The set of such operators is denoted by ${\cal L}_{Leb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{Leb}(X,Y)$).
   2. $c)$
-    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1].
+    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$.
 The set of such operators is denoted by ${\cal L}_{wLeb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{wLeb}(X,Y)$).
   3. $d)$
-    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
 The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf., Definition 2.2d)d)).
-We prove that, under rather mild conditions, quasi $\sigma$-Levi operators are continuous. This result generalizes [11, Theorem 3.1].
+First, we prove that under rather mild conditions, a quasi $\sigma$-Levi operator is continuous.
+This generalizes [11, Theorem 3.1] which states that every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
 Theorem 3.2.
-Every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
-Proof.
-Let $T:E\to F$ be a quasi $\sigma$-Levi operator from a BL $E$ to an NL $F$.
-On the way to a contradiction, suppose $T$ is not continuous.
-Then, there exists a sequence $(x_{n})$ in $B_{X}$ such that $\|Tx_{n}\|\to\infty$.
-Without loss of generality, suppose $(x_{n})\subset(B_{X})^{+}$ and $\|Tx_{n}\|\geq n2^{n}$ for every $n$.
-Set $u=\sum\limits_{k=1}^{\infty}2^{-k}x_{k}$ and $u_{n}=\sum\limits_{k=1}^{n}2^{-k}x_{k}$ for each $n\in\mathbb{N}$.
-Then $0\leq u_{n}\uparrow$ and $\|u_{n}\|\leq 1$ for all $n$.
-Since $T\in\text{\rm L}_{\text{\rm qLevi}}(E,F)$, $(Tu_{n})$ is an o-Cauchy sequence in $F$, and hence the sequence $(Tu_{n})$ is order bounded.
-Thus, $(Tu_{n})$ is also norm bounded.
-It is absurd because $\|Tu_{n+1}-Tu_{n}\|=\|T2^{-(n+1)}x_{n+1}\|\geq n+1$ for all $n$.
-∎
-Theorem 3.3.
-Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered normed space with a normal cone is continuous.
+Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
 Proof.
@@
 As $\{B_{\cal F}(0,2^{-l})\}_{l\in\mathbb{N}}$ is a $\tau$-neighborhood base of zero, the sequence $(u_{i})$ is $\tau$-bounded.
-It follows from $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}({\cal F},Y)$ that $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded.
-Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is also norm bounded.
-It is absurd because
-$\|Tu_{i+1}-Tu_{i}\|=\|Tx_{i+1}\|=\|T(2^{i+1}z_{i+1})\|\geq 2^{i+1}M$
-for all $i\in\mathbb{N}$. ∎
+As $T$ is a quasi $\sigma$-Levi operator, the sequence $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded.
+Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is topologically bounded by [4, Lemma 2.22].
+It is absurd because $Tu_{i+1}-Tu_{i}=Tx_{i+1}=2^{i+1}Tz_{i+1}\notin 2^{i+1}W$ for all $i\in\mathbb{N}$.
+∎
 Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ if and only if $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1].
@@
 Theorem 3.3.
-Let $T$ be a positive operator from an OVS $X$ to an ordered TVS $(Y,\tau)$ with a normal cone. Then
+Let $T$ be a positive operator from an ordered vector space $X$ to an ordered topological vector space $(Y,\tau)$ with a normal cone. Then
 $T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{otc}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{otc}(X,Y).$
@@
 On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{otc}(X,Y)$.
-Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$.
+Then, there exist a net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$.
 Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
 Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$.
-As $T$ is Lebesgue, $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$.
+As $T$ is Lebesgue, $Tg_{\beta}{\xrightarrow[]{\tau}}0$.
 Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$.
@@
 Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$.
-Since $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$.
+Since $Tg_{\beta}{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$.
 Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$.
@@
 Corollary 3.4.
-Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{onc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{onc})_{+}(X,Y)$.
+Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{otc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{otc})_{+}(X,Y)$.
 By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue if and only if $T$ is w-Lebesgue.
@@
 Suppose $T\in{\cal L}_{wLeb}(X,Y)$ and take $x_{\alpha}\downarrow 0$ in $X$.
-Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$.
-By [4, Lemma 2.28], we conclude $Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$.
+Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$.
+By [4, Lemma 2.28], we conclude $Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
 Thus, $T\in{\cal L}_{Leb}(X,Y)$.
@@
 Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{wLeb})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{wLeb})_{+}(X,Y)$.
-Question 3.8.
-Are the conclusions of Corollaries 3.5 and 3.7 hold true for arbitrary (not necessarily positive) operators?
+Question 3.7.
+Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
 The following lemma is a direct consequence of Definition 3.1.
@@
 Lemma 3.9.
-Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\cal L}_{b}(X,Y)$.
+Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then, ${\cal L}_{otb}(X,Y)={\cal L}_{b}(X,Y)$.
 Proof.
-The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
-By [15, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded.
-Accordingly to [15, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous.
-Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
-  1. (*)
-    ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
-Another case of automatic boundedness of operators concerns the order continuity.
-Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have:
-  1. (**)
-    ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
-Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$.
-An inspection of [15, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
+The inclusion ${\cal L}_{otb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
+Another case of automatic continuity of operators concerns their order continuity.
+Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have: ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
+Condition $X_{+}-X_{+}=X$ is essential here since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$.
+Example 3.10.
+([15, Example 2.12 b)]) Choose a normalized Hamel basis $\{\xi:\xi\in\Xi\}$ of an infinite dimensional Banach space $X$ and identify $X$ with $c_{00}(\Xi)$.
+Consider a positive cone $X_{+}:=\Big\{\sum\limits_{\xi\in\Xi}\lambda_{\xi}\xi:\lambda_{\xi}\geq 0\Big\}$ in $X$.
+The cone $X_{+}$ is generating.
+However $X_{+}$ is not closed ((e.g., by [6, Proposition 4.1])).
+Pick any sequence $(\xi_{n})_{n=0}^{\infty}$ of distinct elements of $\Xi$ and define $T:(X,X_{+})\to X$ by $T\xi=0$ for $\xi\in\Xi\setminus\{\xi_{n}:n\in\mathbb{N}\}$ and $T\xi_{n}=n\xi_{0}$ for $n\in\mathbb{N}$.
+It is straightforward to see that the operator $T$ is order-to-norm bounded and positive, but $T$ is not bounded.
 It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [21] without specifying the class of (just linear or continuous linear) operators.
-Automatic boundedness of such operators was investigated in recent paper [22].
-We are going to establish some automatic boundedness conditions for order-to-weak continuous operators on OVSs.
-More precisely, we shall prove Theorem 3.12 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded.
-We begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
+Automatic boundedness of such operators was investigated in the recent paper [22].
+We are going to establish further conditions for automatic boundedness of order-to-weak continuous operators between ordered Banach spaces and and normed spaces.
+More precisely, we prove Theorem 3.12 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded.
+Let us begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
 Lemma 3.11.
-Let $X$ be a normal OBS and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$. In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$.
+Let $X$ be an ordered Banach space with a normal cone and $Y$ a normed space.
+Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$.
+In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$.
 Proof.
@@
 Let $0\leq y_{0}\leq y_{n}$ for all $n\in\mathbb{N}$.
-Since $0\leq y_{0}\leq 2^{1-n}u$ and $\|2^{1-n}u\|\leq M2^{1-n}\to 0$ then $y_{0}=0$ by [4, Theorem 2.23].
+Since $0\leq y_{0}\leq 2^{1-n}u$ it follows that $\|y_{0}\|\leq M2^{1-n}$ for every $n$.
 Thus, $y_{n}\downarrow 0$.
@@
 ∎
-Combining Lemmas 3.9, 3.10, and 3.11, we obtain.
+Combining Lemmas 3.8, 3.9, and 3.11, we obtain.
 Theorem 3.12.
-Let $X$ be an ordered Banach space with a closed generating normal cone and let $Y$ be a normed space.
+Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space.
 Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
@@
 Corollary 3.13.
-(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
-The next result is a straightforward consequence of Lemma 3.9 and Theorem 3.12.
-Corollary 3.14.
-Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS.
-Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\cal L}_{b}(X,Y)$.
-Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1.
-c)]).
+(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an normed space is bounded.
+Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded [5, Remark 1.1.
+c)].
 From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1].
 Definition 3.14.
-The topology in an ordered TVS $(X,\tau)$ (the norm in an ONS $X$) is
-  1. $a)$
-    order continuous if $x_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{\alpha}\|\to 0$) for every net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-  2. $b)$
-    $\sigma$-order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+The norm in an ordered normed space $X$ is order continuous if $\|x_{\alpha}\|\to 0$ for every net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
 In certain cases o-convergence is equivalent to ru-convergence.
 Lemma 3.15.
-Let $X$ be an OBS with a closed normal cone. The following assertions hold.
-  1. $i)$
-    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
-  2. ii)ii)
-    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
+Let $X$ be an OBS with a closed normal cone and the norm in $X$ is order continuous. Then, $x_{\alpha}{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}{\xrightarrow[]{ru}}0$ in $X$.
 Proof.
-First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-Suppose $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
-Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$ with $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
-Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$.
-Since $\big\|\frac{1}{n}u\big\|\to 0$ then $y_{0}=0$ by [4, Theorem 2.23].
-It follows $\frac{1}{n}u\downarrow 0$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$.
-$i)$ Let $X$ has order continuous norm and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+Since $X_{+}$ is normal, $X$ is Archimedean, and hence $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
+Let $X$ has order continuous norm and $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
 Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
-Find an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$, and set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$.
+Since the norm in $X$ is order continuous and $X_{+}$ is normal, there exists an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$.
+Set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$.
 Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$.
 Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$.
-Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
-ii)ii) The proof is similar. ∎
+Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}{\xrightarrow[]{ru}}0$.
+∎
 Theorem 3.16.
-Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. The following assertions hold.
-  1. $i)$
-    If $X$ has $\sigma$-order continuous norm then L^σ_Leb(X,Y)=L^σ_onc(X,Y) and L^σ_wLeb(X,Y)=L^σ_owc(X,Y).
-  2. ii)ii)
-    If $X$ has order continuous norm then L_Leb(X,Y)=L_onc(X,Y) and L_wLeb(X,Y)=L_owc(X,Y).
+Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. If $X$ has order continuous norm then
+${\cal L}_{Leb}(X,Y)={\cal L}_{otc}(X,Y)\ \ and\ \ {\cal L}_{wLeb}(X,Y)={\cal L}_{owc}(X,Y).$
 Proof.
-We prove only $i)$ as the proof of ii)ii) is similar.
-In view of Lemma 3.9, we need to prove ${\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$ and ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y)$.
-Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-Then $x_{n}\operatorname{\xrightarrow[]{ru}}0$, by Lemma 3.16, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
-If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12, and hence $\|Tx_{n}\|\to 0$.
-If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$.
-Then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12.
-So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$.
+In view of Lemma 3.8, we need to prove ${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{otc}(X,Y)$ and ${\cal L}_{wLeb}(X,Y)\subseteq{\cal L}_{owc}(X,Y)$.
+Let $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
+Then $x_{\alpha}{\xrightarrow[]{ru}}0$ by Lemma 3.15, and hence $\|x_{\alpha}\|\to 0$ due to normality of $X_{+}$.
+If $T\in{\cal L}_{Leb}(X,Y)$ then $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$.
+By Theorem 3.12, $T\in{\cal L}_{b}(X,Y)$, and hence $\|Tx_{\alpha}\|\to 0$.
+Thus, $T\in{\cal L}_{otc}(X,Y)$.
+If $T\in{\cal L}_{wLeb}(X,Y)$ then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12.
+Thus, $\|Tx_{\alpha}\|\to 0$, and hence $Tx_{\alpha}{\xrightarrow[]{w}}0$.
+So, $T\in{\cal L}_{owc}(X,Y)$.
 ∎
@@
 Corollary 3.17.
-Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space.
-Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
-Finally, we give some conditions on order bounded operators to be order-to-norm continuous.
+Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space. Then, ${\cal L}_{Leb}(E,Y)={\cal L}_{otc}(E,Y)$ and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
+Finally, we give the conditions for order-to-topology continuity of order bounded operators.
 Theorem 3.18.
-Let $X$ be an ordered Banach space with a closed normal cone and $Y$ an ordered normed space with a generating normal cone. The following assertions hold.
-  1. $i)$
-    If $X$ has $\sigma$-order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$.
-  2. ii)ii)
-    If $X$ has order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{onc}(X,Y)$.
+Let $X$ be an ordered Banach space with a closed normal cone and order continuous norm, $Y$ be an ordered normed space with a generating normal cone, and $T:X\to Y$ and order bounded operator.
+Then $T$ is order-to-topology continuous.
 Proof.
-We prove only $i)$ as the proof of ii)ii) is similar.
-Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
-By Lemma 3.16, $x_{n}\operatorname{\xrightarrow[]{ru}}0$.
-It follows from [15, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{ru}}0$.
-The normality of the cone $Y_{+}$ ensures $\|Tx_{n}\|\to 0$.
+Let $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
+By Lemma 3.15, $x_{\alpha}{\xrightarrow[]{ru}}0$.
+It follows from [15, Corollary 2.5] that $Tx_{\alpha}{\xrightarrow[]{ru}}0$.
+The normality of the cone $Y_{+}$ ensures $\|Tx_{\alpha}\|\to 0$.
 ∎